As a finance manager of Kersot Enterprise Limited, where the Board of Trustees is reviewing the mix of the capital structure of the enterprise based on target value and has decided to raise 40 percent of new funds from long-term debt, 10 percent from preferred shares, and 50 percent from ordinary shares with components of 5.6, 10.5 and 15.7 respectively.

Estimate percentage of this BEST MIX for the enterprise and support the decision taken.           


1
Expert's answer
2022-02-21T13:03:55-0500

Capital Structure proportion is as follows:

Proportion of long term debt=40%=0.

Proportion of ordinary shares=50%=0.5

Proportion of Preferred share=10%=0.1

weighted average cost(WACC)

WACC =("K_{e}" "\times" "P_{e}" )+("K_{d}" "\times" "P_{d}" )+("K_{pa}" "\times" "P_{pa}" )

WACC =0.4"\times" 5.6+ 0.1"\times" 10.5+0.5"\times" 15.7=11.14%

The manager of Kersot Enterprise Limited should ;

 lower the cost of equity or change the capital structure to include more debt. Since the cost of equity reflects the risk associated with generating future net cash flow, lowering the company's risk characteristics will also lower this cost.